OpenStudy (anonymous):

find the least common denominator of 3/4 & 5/6

OpenStudy (anonymous):

4*2 = 8. Does 6 go into 8? Nope. 4*3 = 12 Does 6 go into 12? Awww yeah. 6*2 = 12.

OpenStudy (anonymous):

6*2 = 12. 12 divisible by 9? No. 6*3 = 18. 18 divisible by 9? You tell me.

OpenStudy (anonymous):

i thoght the least common denominator was 60 and i dont understand why it isint

OpenStudy (anonymous):

hello???